### 15 Apr 2026 v2

# Genus three Ceresa cycles and limit of archimedean heights

Abstract.
For a one-parameter variation of biextension mixed Hodge structures, Brosnan and Pearlstein showed that the limit of the asymptotic height of the variation is given by a certain limit height of the nilpotent orbit. This limit height depends on the choice of a parameter. In the case of a variation of geometric origin related to Ceresa cycles associated with curves of genus three, after fixing a parameter, we show that this limit height is given by the Deligne splitting of a biextension mixed Hodge structure associated with cycles in the boundary.
2020 Mathematics Subject Classification:
14C25, 14C30, 37P30
S. Goswami was partially supported by the Universitat de Barcelona María Zambrano postdoctoral fellowship. I. Spelta is a member of GNSAGA (INdAM). Both authors were partially supported by the Spanish MINECO research project PID2023-147642NB-I00.
Introduction
The pair $(C,p)$, consisting of a smooth projective complex algebraic curve $C$ together with a marked point $p$, determines an algebraic $1$-cycle on its Jacobian $JC$, known as the Ceresa cycle, defined by $Z(C):=C-[-1]^{*}C.$ By construction, the Ceresa cycle is homologically trivial. However, a fundamental result of Ceresa ([Ce83]) shows that for a general curve of genus $g\geq 3$, it is not algebraically equivalent to zero. This nontriviality has made the Ceresa cycle a central object in the study of the geometry of curves and their Jacobians.
Let $\Delta\coloneqq\{t\in\mathds{C}:\penalty 10000\ |t|<1\}$ be the open unit disc. Consider a family $\mathcal{C}\to\Delta$ of curves of fixed genus $g\geq 3$, with smooth and projective general fiber $C_{t},\penalty 10000\ t\neq 0$, and central fiber $C_{0}$ having nodal singularities. Two functions can naturally be attached to such a family. The first one is a holomorphic function $\nu(t)\coloneqq AJ(Z(C_{t}))$, known as a normal function and defined using the Abel-Jacobi image of the Ceresa cycle. Both $\nu$ and its derivative, the Griffiths infinitesimal invariant, have been studied widely by several authors (see, for example, [Gr83], [Voi88], [Ha90], [CP95],[Plz03], [GGK10], and [Ha25]). The second function is the height, which we introduce more generally. Let $X$ be a smooth complex projective variety, and let $Z$ and $W$ be homologically trivial algebraic cycles of codimensions $p$ and $q$, respectively, intersecting properly and satisfying $\dim X+1=p+q$. Under these assumptions, one necessarily has $|Z|\cap|W|=\emptyset$. To such a pair, one can associate a unique real number $\operatorname{ht}(Z,W)$, called the archimedean height. Hodge theoretically, $\operatorname{ht}(Z,W)$ measures how far a canonical subquotient $B_{Z,W}$ of the (rational) mixed Hodge structure $H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))$ is from being $\mathds{R}$-split. $B_{Z,W}$ is an example of a biextension mixed Hodge structure with graded pieces
$\operatorname{Gr}^{W}_{0}B_{Z,W}=\mathds{Q}(0),\penalty 10000\ \operatorname{Gr}^{W}_{-1}B_{Z,W}=H^{2p-1}(X;\mathds{Q}(p)),\penalty 10000\ \operatorname{Gr}^{W}_{-2}B_{Z,W}=\mathds{Q}(1),$
and $\operatorname{Gr}^{W}_{k}B_{Z,W}=0$ otherwise. Conceptually, we can say that $B_{Z,W}$ ‘splices’ together the Abel-Jacobi class $AJ(Z)$ and the dual $AJ(W)^{\vee}$ of the Abel-Jacobi class of $W$. One can define a metric on $B_{Z,W}$ by $B_{Z,W}||\coloneqq e^{\operatorname{ht}(Z,W)}$. Similarly, for an abstract biextension $B$ (see section 3 for the definition) one can attach a height $\operatorname{ht}(B)$, and define a metric $B||=e^{\operatorname{ht}(B)}$.
Using the above definition of metric on biextensions and Ceresa cycles, in [HR04], Hain and Reed introduced a naturally metrized line bundle $(\mathcal{B},\|\cdot\|_{\mathcal{B}})$ on the moduli space $\mathcal{M}_{g}$ of smooth projective curves of genus $g\geq 3$, called the biextension line bundle. The associated metric encodes the archimedean contribution of the Beilinson height of pairs of Ceresa cycles.
A natural problem is to extend this construction from $\mathcal{M}_{g}$ to its Deligne–Mumford compactification $\overline{\mathcal{M}}_{g}$. It can be shown that while the biextension line bundle extends algebraically, the metric does not (see [Lear90], [HR04], [Pe06, Section 5]). In the language of heights, for a family of biextensions $\{B_{t}\}_{t\neq 0}$, instead of $\operatorname{ht}(B_{t})$, the asymptotic function $\operatorname{ht}(B_{t})+\mu\log|t|$ extends continuously to $t=0$, for $\mu\in\mathds{Q}$ that is controlled by the monodromy around $t=0$. This $\mu$ measures the singular behaviour of the height near the boundary. Then, the above study was generalized for an abstract biextension variation over a more general base in [BP19].
While the focus of the existing literature has been the study of $\mu$ and, more generally, the height jump divisor, the aim of our paper is slightly different. In [BP19, Definition 70], the authors introduce a limit height $H(N,F_{\infty},W)$ of the nilpotent orbit by ([BP19, Theorem 76])
$\lim_{t\to 0}\Big(\operatorname{ht}(B_{t})+\mu\log|t|\Big)=H(N,F_{\infty},W).$
Much like the limit mixed Hodge structure, the limit height depends on the choice of coordinate, while $\mu$ is independent of it. In this article, we aim to show that in a biextension variation of geometric origin, after fixing a choice of coordinate for $\Delta$, $H(N,F_{\infty},W)$ is ‘given by a biextension mixed Hodge structure attached to cycles in the boundary’. More precisely, following the approaches of [Co97, Ha90, Plz03], we consider the limit of asymptotic heights of Ceresa cycles associated to a special family of genus three curves as the first non-trivial case. We summarize our results as follows.
Main Theorem.
Let $\mathcal{C}\to\Delta$ be a family of genus $g=3$ curves whose fiber $C_{t}$ is smooth and projective for $t\neq 0$, and whose central fiber $C_{0}$ is one of the following two types:
  • $C_{0}$ is an irreducible nodal curve with a single node $x_{0}$, such that for a normalization $(\widetilde{C}_{0},\{p,q\})\to(C_{0},\{x_{0}\})$, the zero cycle pqp-q is torsion in $J\widetilde{C}_{0}$ (assumption 1.8).
  • $C_{0}=D\cup E$, for two irreducible and smooth components $E$ and $D$ of genus one and two, respectively, with a separating node $y_{0}=D\cap E$.
Then the following happens:
  1. (i)
    When $C_{0}$ is irreducible and nodal, assuming 1.8, there is a choice of coordinate (assumption 2.2) such that the associated limit mixed Hodge structure $H^{1}_{\lim}(C_{t};\mathds{Q})$ is split. When $C_{0}$ is reducible, $H^{1}_{\lim}(C_{t};\mathds{Q})\cong H^{1}(D;\mathds{Q})\oplus H^{1}(E;\mathds{Q})$ is automatically split.
  2. (ii)
    There exists a family of Ceresa cycles $\{Z(C_{t}),W(C_{t})\}_{t\in\Delta^{\ast}}$ that intersects properly in $JC_{t}$, such that the cycles $Z(C_{0})$ and $W(C_{0})$ also intersect properly in the compactification $X_{0}$ of the Jacobian $JC_{0}$, and are homologous to zero (Propositions 4.1 and 4.5).
  3. (iii)
    Under assumptions 1.8 and 2.2, the limit height $H(N,F_{\infty},W)$ associated with the biextension variation $B_{Z(C_{t}),W(C_{t})}$, is given by the Deligne splitting of the biextension mixed Hodge structure $B_{\widetilde{Z}(C_{0}),\widetilde{W}(C_{0})}$, where $\widetilde{Z}(C_{0})$ and $\widetilde{W}(C_{0})$ are the lifts of $Z(C_{0})$ and $W(C_{0})$ to the normalization $\widetilde{X}_{0}$ and are homologically trivial.
We mention that if $C_{0}$ is nodal and irreducible, there is a nontrivial vanishing cycle and a monodromy operator $N\neq 0$ with $N^{2}=0$ for $H^{1}(C_{t};\mathds{Q})$. The case where $C_{0}$ is reducible is simpler, since $N=0$ on $H^{1}(C_{t};\mathds{Q})$. Thus, in this case, the choice of coordinate is not explicit. Moreover, $JC_{0}$ is smooth and compact, so no normalization is needed in $(iii)$.
We now give an idea of the proof when $C_{0}$ is nodal and irreducible. Our method is inspired by the work in [Plz03] and [BP19], and can be thought of as a bridge between the two. Using assumption 1.8, we can fix a coordinate in such a way that the mixed Hodge structure $H^{1}_{\lim}(C_{t};\mathds{Q})$ is split (assumption 2.2). Then the key step involves finding the limit of the biextension variation
$\mathcal{V}_{t}=H^{3}(JC_{t}\setminus|Z(C_{t})|,|W(C_{t})|;\mathbb{Q}(2)).$
The assumption 2.2 makes this study significantly easier. Nevertheless, the resulting limit mixed Hodge structure $\mathcal{V}_{\lim}$ no longer has the structure of a biextension. Consequently, the analysis of the limiting height amounts to extracting from $\mathcal{V}_{\lim}$ the graded pieces of the biextension $H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}(C_{0})|,|\widetilde{W}(C_{0})|;\mathds{Q}(2))$. This is the core motivation of our result (see Theorem 4.7), and is achieved using a purity result for the cohomology of ‘mildly’ singular varieties (see Proposition 4.3), and the injectivity of the Clemens-Schmidt morphism in our situation. Finally, we conclude by exploiting a relation between heights and certain morphisms of mixed Hodge structures. (see Proposition 3.4).
Remark.
Main Theorem(iii) is not merely an equality of two real numbers. The main point is that the limit height $H(N,F_{\infty},W)$ is obtained as a Deligne splitting of a geometrically defined biextension mixed Hodge structure appearing in the boundary. Heuristically, we are studying the interaction between the two mixed Hodge structures $H^{3}_{\lim}(JC_{t}\setminus|Z(C_{t})|,|W(C_{t})|;\mathds{Q}(2))$ and $H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}(C_{0})|,|\widetilde{W}(C_{0})|;\mathds{Q}(2))$, which is much deeper than simply the equality of two invariants as real numbers.
The paper is organized as follows. In Section 1, we review basic material on mixed Hodge structures and on the Jacobian of a uni-nodal curve, and we introduce the two types of degenerations used throughout the paper. In Section 2, we recall the theory of limit mixed Hodge structures of geometric origin and apply it to analyze the mixed Hodge structure $H^{3}_{\lim}(JC_{t};\mathds{Q})$. In Section 3, we discuss biextensions, both abstractly and in the geometric setting, introduce the notion of height, and recall the variational aspects of [BP19]. Finally, in Section 4, we study families of Ceresa cycles, establish their properties and prove our main theorem (Theorem 4.7).
1. Preliminaries
In this section, we gather preliminary results from Hodge theory, nodal curves, and Jacobians.
1.1. Mixed Hodge structures
In this section, we introduce rational mixed Hodge structures and list some properties (mostly without proofs) that we will need. We point out to the reader that working with rational mixed Hodge structures is a matter of convenience, more than anything else. One may as well work with integral structures modulo torsion, and the results of this article will remain valid.
Definition 1.1.
A rational mixed Hodge structure is a triple
$(H,W_{\bullet},F^{\bullet})$
consisting of a $\mathds{Q}$-vector space $H$, a finite increasing filtration $W_{\bullet}$ of $H$ (called the weight filtration) and a finite decreasing filtration $F^{\bullet}$ of $H_{\mathds{C}}=H\otimes\mathds{C}$ (called the Hodge filtration) such that for each $n\in\mathds{Z}$ the couple $(\operatorname{Gr}_{n}^{W}H,\operatorname{Gr}_{n}^{W}(F^{\bullet}))$ is a pure $\mathds{Q}$-Hodge structure of weight $n$. One can also consider real mixed Hodge structures, where instead of a $\mathds{Q}$-vector space $H_{\mathds{Q}}$ one has a real vector space $H_{\mathds{R}}$. In fact, given a mixed $\mathds{Q}$-Hodge structure $H$, we will denote $H_{\mathds{R}}=H_{\mathds{Q}}\otimes\mathds{R}$ obtaining an $\mathds{R}$-mixed Hodge structure.
When studying variations of mixed Hodge structures, it is convenient to fix the underlying vector space and move the filtrations $F$ and $W$. Thus if we often fix an ($\mathds{R}$ or $\mathds{Q}$) vector space $V$, then a pair of filtrations $(F,W)$ on $V\otimes\mathds{C}$ and $V$ respectively is called a mixed Hodge structure if the triple
$((V,W),(V\otimes\mathds{C},W,F),\operatorname{Id}_{V\otimes\mathds{C}})$
is a mixed Hodge structure.
Example 1.2.
For $a\in\mathds{Z}$, the Tate mixed Hodge structure $\mathds{Q}(a)$ is the mixed Hodge structure given by the following data
$\mathds{Q}(a)_{\mathds{Q}}=\mathds{Q},\quad W_{-2a-1}\mathds{Q}(a)_{\mathds{Q}}=0,\quad W_{-2a}\mathds{Q}(a)_{\mathds{Q}}=\mathds{Q}$
$\mathds{Q}(a)_{\mathds{C}}=\mathds{C},\quad F^{-a}\mathds{Q}(a)_{\mathds{C}}=\mathds{C},\quad F^{-a+1}\mathds{Q}(a)_{\mathds{C}}=0$
Definition 1.3.
A morphism $f\colon(H,F,W)\to(H^{\prime},F^{\prime},W^{\prime})$ of mixed Hodge structures is a $\mathds{Q}$-linear map which is compatible with the Hodge and weight filtrations $(F,W)$ and $(F^{\prime},W^{\prime})$ respectively. For ease of notation, we will omit the Hodge and weight filtrations and denote a morphism of mixed Hodge structures simply by $f\colon H\to H^{\prime}$.
1.1.1. Deligne splitting
A mixed Hodge structure $(F,W)$ on $V$ induces a unique functorial bigrading
(1.1) $V_{\mathds{C}}=\bigoplus_{a,b}\,I^{a,b}$
of the underlying complex vector space $V_{\mathds{C}}$ such that
  1. (i)
    $F^{a}=\oplus_{\alpha\geq a,\beta}\,I^{\alpha,\beta}$;
  2. (ii)
    $W_{k}=\oplus_{\alpha+\beta\leq k}\,I^{\alpha,\beta}$;
  3. (iii)
    $\overline{I^{a,b}}\equiv I^{b,a}\mod\oplus_{\beta<b,\alpha<a}\,I^{\beta,\alpha}$.
The $I^{a,b}$ is given by
$I^{a,b}=F^{a}\cap W_{a+b}\cap\left(\overline{F^{b}}\cap W_{a+b}+\overline{U^{b-1}_{a+b-2}}\right),$
where
$U^{r}_{s}=\sum_{j\geq 0}F^{r-j}\cap W_{s-j}.$
A mixed Hodge structure is said to be split (over $\mathds{R}$) if $\bar{I}^{a,b}=I^{b,a}$.
Definition 1.4.
The bigrading (1.1) will be called the Deligne bigrading of $(F,W)$. The associated semisimple endomorphism $Y=Y_{(F,W)}$ of $V_{\mathds{C}}$ which acts as multiplication by $p+q$ on $I^{p,q}$ will be called the Deligne grading of $(F,W)$.
An important consequence of the Deligne bigrading is the strictness of a morphism of mixed Hodge structures with respect to the Hodge and the weight filtrations. Moreover, one also has
Corollary 1.5.
A short exact sequence
$0\to H^{{}^{\prime}}\to H\to H^{{}^{\prime\prime}}\to 0$
induces a short exact sequence on graded pieces
$0\to\operatorname{Gr}^{W}H^{{}^{\prime}}\to\operatorname{Gr}^{W}H\to\operatorname{Gr}^{W}H^{{}^{\prime\prime}}\to 0.$
Following Cattani, Kaplan, and Schmid ([CKS86]), let
$\operatorname{\mathfrak{gl}}(V_{\mathds{C}})^{a,b}=\{\,\alpha\in\operatorname{\mathfrak{gl}}(V_{\mathds{C}})\mid\alpha(I^{c,d})\subseteq I^{a+c,b+d}\,\}$
be the Hodge decomposition of $\operatorname{\mathfrak{gl}}(V_{\mathds{C}})$: an element $\lambda\in\operatorname{\mathfrak{gl}}(V_{\mathds{C}})$ decomposes as $\lambda=\sum\lambda^{a,b}$. Defining
(1.2) $\Lambda^{-1,-1}=\bigoplus_{a<0,b<0}\,\operatorname{\mathfrak{gl}}(V_{\mathds{C}})^{a,b},$
then we have $\overline{\Lambda^{-1,-1}}=\Lambda^{-1,-1}$. It follows from the properties of the Deligne bigrading that taking $\lambda\in\Lambda^{-1,-1}$ then $(e^{\lambda}\cdot F,W)$ is another mixed Hodge structure on $V$ such that
[…]
In particular, there exists a unique real element $\delta=\delta_{(F,W)}\in\Lambda^{-1,-1}$ such that
(1.3) $\overline{Y}_{(F,W)}=e^{-2i\delta}\cdot Y_{(F,W)}$
where $g\cdot\alpha\coloneqq\operatorname{Ad}(g)\alpha$ denotes the adjoint action of $\operatorname{GL}(V_{\mathds{C}})$ on $\operatorname{\mathfrak{gl}}(V_{\mathds{C}})$. The element $\delta$ defined by (1.3) will be called the Deligne delta of $(F,W)$. It is responsible for turning a non-split mixed Hodge structure into a split one.
Remark 1.6.
Let $\delta_{-j}\colon\operatorname{Gr}^{W}_{i}V_{\mathds{C}}\to\operatorname{Gr}^{W}_{i-j}V_{\mathds{C}}$ be the component of $\delta$ that maps the $i$-th graded piece to the iji-j-th graded piece, for all $i\in\mathds{Z}$. Then
$\text{ad}(Y)(\delta_{-j})(v)=(i-j)\delta_{-j}(v)-i\delta_{-j}(v)=-j\delta_{-j}(v).$
So $\delta_{-j}$ belongs to the $j-$th eigenspace of $\text{ad}(Y)$, and we have a decomposition $\delta=\Sigma_{j}\delta_{-j}$ of $\delta$ in terms of these eigencomponents. Notice that, since $\delta\in\Lambda^{-1,-1}$, by (1.2), $\delta_{-j}=\sum_{a+b=-j}\delta_{-j}^{a,b}$ where $a<0$ and $b<0$. Thus, the smallest possible shift is $j=2$.
Next, we show that $\delta$ commutes with morphisms of mixed Hodge structures.
Lemma 1.7.
Let $A$ and $B$ be mixed Hodge structures with Deligne splittings $\delta_{A}$ and $\delta_{B}$ respectively. Let $f:A\to B$ be a morphism of mixed Hodge structures. Then, $f\circ\delta_{A}=\delta_{B}\circ f$.
Proof.
By [CKS86, Proposition 2.20], if $C$ is a mixed Hodge structure, then $\delta_{C}$ commutes with all $(r,r)$-morphisms of $C$. Let $C=A\oplus B$ and observe that $g(a,b)=(a,b+f(a))$ is a morphism of $C$. Using the block structure of $\operatorname{\mathfrak{gl}}(C)=\operatorname{\mathfrak{gl}}(A\oplus B)$ it follows immediately from (1.3) that $\delta_{C}(a,b)=(\delta_{A}(a),\delta_{B}(b))$. Writing out the $g\circ\delta_{C}=\delta_{C}\circ g$ shows that $f\circ\delta_{A}=\delta_{B}\circ f$. ∎
1.2. Nodal curves and Jacobians
Next, we survey the structure of the Jacobian of a nodal curve and discuss the Abel Jacobi maps.
1.2.1. Jacobian of a nodal curve
Let $C_{0}$ be a connected irreducible nodal curve of genus $g$ with one node (ordinary double point) $x_{0}$ as a singularity. We define $JC_{0}$ as the variety of isomorphism classes of line bundles of degree zero on $C_{0}$, and the tensor product yields on $JC_{0}$ the structure of a group variety. Let $\pi\colon(\widetilde{C}_{0},\{p,q\})\rightarrow(C_{0},\{x_{0}\})$ be a normalization of $C_{0}$, where $\widetilde{C}_{0}$ is smooth projective curve of genus g1g-1 (opening the node), such that $\pi$ defines an isomorphism $\widetilde{C}_{0}\setminus\{p,q\}\cong C_{0}\setminus\{x_{0}\}$. Then, we define the pullback morphism
$\pi^{\ast}\colon JC_{0}\rightarrow J\widetilde{C}_{0}={\rm Pic}^{0}(\widetilde{C}_{0}).$
As it is well-known, this morphism is surjective, and the kernel is given by $\mathds{C}^{\ast}$. Hence, we have an exact sequence which presents $JC_{0}$ as a semi-abelian variety:
$0\rightarrow\mathds{C}^{\ast}\rightarrow JC_{0}\xrightarrow{\pi^{\ast}}J\widetilde{C}_{0}\rightarrow 0.$
The semi-abelian variety $JC_{0}$ can be compactified. Indeed, identifying ${\rm Pic}^{0}(\widetilde{C}_{0})\cong{\rm Pic}^{0}(J\widetilde{C}_{0})$, the line bundle $L:=\mathcal{O}_{\widetilde{C}_{0}}(p-q)$ can be viewed as a line bundle over $J\widetilde{C}_{0}$, and $JC_{0}$ is the total space of $L$ minus the 0-section. Following [AK90], we give a quick description of the compactification: Let $Y_{0}$ and $Y_{\infty}$ be the zero and infinity sections of $\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L)$ obtained through the projections $s_{0}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to\mathcal{O}_{J\widetilde{C}_{0}}$ and $s_{1}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to L$ respectively. Then, the compactification $X_{0}=\overline{JC_{0}}$ is obtained as
$X_{0}=\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L)/Y_{0}\sim Y_{\infty}.$
Notice that the two sections $Y_{0}$ and $Y_{\infty}$ do not get identified fiberwise. Indeed, $L$ being an element in $J\widetilde{C}_{0}$, the identification works as
(1.4) $(D,0)\in Y_{0}\sim(D+p-q,\infty)\in Y_{\infty}.$
As sets we have:
$X_{0}:=\overline{JC_{0}}=JC_{0}\coprod J\widetilde{C}_{0},$
in other words the singular locus of $X_{0}$ is given by $J\widetilde{C}_{0}$, which is smooth and projective. Elements in $X_{0}$ are rank one torsion-free sheaves $M$ of degree 0 whose sections $s$ are allowed to approach the point $x_{0}$. The normalization $\widetilde{X}_{0}$ is given by the projective bundle $\mathds{P}\left(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\right)$ over $J\widetilde{C}_{0}$. From [UP24, Proposition 5.3] and the projective bundle formula for cohomology, the singular cohomology groups of $\widetilde{X}_{0}$ and $X_{0}$ decompose:
(1.5) $H^{k}(\widetilde{X}_{0};\mathds{Q})\cong H^{k}(J\widetilde{C}_{0};\mathds{Q})\oplus H^{k-2}(J\widetilde{C}_{0};\mathds{Q}(-1)).$
and
(1.6) $H^{k}(X_{0};\mathds{Q})\cong H^{k}(\widetilde{X}_{0};\mathds{Q})\oplus H^{k-1}(J\widetilde{C}_{0};\mathds{Q}),$
Notice that both are isomorphisms of mixed Hodge structures. The weight $k$ component of $H^{k}(X_{0};\mathds{Q})$ is given by $H^{k}(\widetilde{X}_{0};\mathds{Q})$, and the k1k-1 component is given by $H^{k-1}(J\widetilde{C}_{0};\mathds{Q})$. However, the decomposition (1.6) is non-canonical. On the other hand, since $\widetilde{X}_{0}$ is smooth and projective, $H^{k}(\widetilde{X}_{0};\mathds{Q})$ carries a pure Hodge structure.
1.2.2. Abel Jacobi map on stable curves
It is well-known that, when $C$ is a smooth projective genus $g$ curve, fixing a base point $x$, the Abel Jacobi map is the embedding:
$AJ:C\hookrightarrow JC,c\mapsto\mathcal{O}_{C}(c-x).$
In the last few decades, several authors have considered the problem of defining a similar map in the case of singular curves. The following is borrowed from [CCE08]. Let $C_{0}$ be a reduced, reducible, connected curve, and for each point $c\in C_{0}$, let $\mathcal{I}_{c}$ be its sheaf of ideals, which is torsion-free, rank 1, and of degree -1, and simple if $c\in C_{0}$ is a non-separating node. Then, if $C_{0}$ is free from separating nodes, and $x\in C_{0}$ is a smooth point,
(1.7) $AJ:C_{0}\rightarrow X_{0},\;c\mapsto\mathcal{I}_{c}\otimes\mathcal{O}_{C_{0}}(x)$
is the degree 0 Abel map of $C_{0}$ with base point $x$. Furthermore, if $C_{0}\neq\mathbb{P}^{1}$, then (1.7) is an embedding. When $c$ is a smooth point, $\mathcal{I}_{c}=\mathcal{O}_{C_{0}}(-c)$, but, if $c$ is a node, then $\mathcal{I}_{c}$ is no longer locally free, and thus $AJ(c)$ is a boundary point in $\partial X_{0}$.
On the other hand, if $C_{0}$ has a separating node $y_{0}$, the map (1.7) has to be corrected, since multidegrees appear. In this case, $X_{0}$ parametrizes isomorphism classes of line bundles having degree 0 on every irreducible component. Furthermore, for each tail $Z$ of $C_{0}$, we consider the line bundles $\mathcal{O}_{C_{0}}(Z)$ which restricts to $Z$ as $\mathcal{O}_{Z}(-y_{0})$, while on the complementary curve $Z^{\prime}$ as $\mathcal{O}_{Z^{\prime}}(y_{0})$. These are usually called twisters. Then for $c\in C_{0}$ such that $c\neq y_{0}$, the sheaf $\mathcal{I}_{c}$ is, as before, the sheaf of ideals of $c$, while $\mathcal{I}_{y_{0}}$ is the unique (simple) line bundle such that
$\mathcal{I}_{y_{0}}|_{Z}\cong\mathcal{O}_{Z}(-y_{0})\quad\mathcal{I}_{c}|_{Z^{\prime}}\cong\mathcal{O}_{Z^{\prime}},$
where $Z$ is the $x$-tail not containing $x$, and $Z^{\prime}$ the complementary. Hence, the degree 0 twisted Abel map of $C_{0}$ with base point $x$ is defined as
(1.8) $AJ:C\rightarrow X_{0},\;c\mapsto\mathcal{I}_{c}\otimes\mathcal{O}_{C_{0}}(x)\otimes\mathcal{O}_{C_{0}}(-\sum_{Z\in\mathcal{T}_{x}(X)}Z),$
where $\mathcal{T}_{x}(X)$ is the set of all $x$-tails.
1.2.3. Degenerations
As is customary, let $\overline{\mathcal{M}}_{g}$ be the moduli space of stable genus $g$ algebraic curves. The complement $\overline{\mathcal{M}}_{g}\smallsetminus\mathcal{M}_{g}$ is a divisor $D$ which is the union of $[g/2]+1$ irreducible components $D=D_{0}\cup D_{1}\cup\ldots\cup D_{[g/2]}$. The general point of $D_{0}$ is an irreducible stable curve of genus $g$ with one node, while the general point of $D_{i}$ for $i>0$ consists of a smooth curve of genus $i$ joined at one point to a smooth curve of genus gig-i. In this article, we will focus on $g=3$ and look at the general element in the boundary divisors. The central fibre $C_{0}$ of our family will be of the following two types.
  • 1.
    An irreducible curve $C_{0}$ with one node at $x_{0}\in C_{0}$. Moreover, we work under the following assumption.
    Assumption 1.8.
    Let $(\widetilde{C}_{0},\{p,q\})\to(C_{0},\{x_{0}\})$ be a normalization of $C_{0}$. We assume that the zero cycle pqp-q is torsion in $J\widetilde{C}_{0}$.
    We remark here that even though this torsion condition is not very natural in general, the example we have in the back of our mind is the degeneration of a family of Ceresa cycles associated to a family of genus three curves, to the Collino cycle, as constructed in [Co97, Section (1.2)].
  • 2.
    A reducible curve $C_{0}=C_{1}\cup C_{2}$ obtained as the union of two irreducible smooth components $C_{j}$ of genus $1$ and $2$ respectively, with a separating node $\{y_{0}\}=C_{1}\cap C_{2}$. Let $q_{i}\in C_{i},i=1,2$ lying above $y_{0}$. In this case $X_{0}\cong JC_{1}\times JC_{2}.$
In particular, (1.7) can be seen as $AJ:C_{0}\smallsetminus x_{0}\rightarrow JC_{0},\;c\mapsto\mathcal{O}_{C_{0}}(c-x)$ extended to the node via the compactification $X_{0}$. On the other hand, (1.8) works as follows:
$q\in C_{1}$ $\mapsto(\mathcal{O}_{C_{1}}(q-q_{1}),\mathcal{O}_{C_{2}}(q_{2}-x))$
(1.9) $q\in C_{2}$ $\mapsto(\mathcal{O}_{C_{1}},\mathcal{O}_{C_{2}}(q-x))$
$y_{0}$ $\mapsto(\mathcal{O}_{C_{1}},\mathcal{O}_{C_{2}}(q_{2}-x)).$
Thus, the image of $C_{1}$ in $JC_{1}$ is the Abel Jacobi curve with base point $q_{1}$, while the image of $C_{2}$ in $JC_{2}$ is the Abel Jacobi curve with base point $x$. Notice that, when the base point $x$ belongs to $C_{1}$, (1.8) reads as:
$q\in C_{1}$ $\mapsto(\mathcal{O}_{C_{1}}(x-q),\mathcal{O}_{C_{2}})$
(1.10) $q\in C_{2}$ $\mapsto(\mathcal{O}_{C_{1}}(x-q_{1}),\mathcal{O}_{C_{2}}(q-q_{2}))$
$y_{0}$ $\mapsto(\mathcal{O}_{C_{1}}(x-q_{1}),\mathcal{O}_{C_{2}}).$
2. Degeneration of curves and their Jacobians
In this section, we discuss degenerations of mixed Hodge structures of geometric origin. In particular, we focus on $H^{1}(C_{t};\mathds{Q})$ for families of smooth projective curves $C_{t}$, and we discuss the mixed Hodge structure we obtain in the limit of our families. We fix the genus at $g=3$ to maintain parity with our main discussion, noting that similar results hold for general genera.
2.1. Survey of general degenerations of geometric origin
Let us set $\Delta\coloneqq\{z\in\mathds{C};\penalty 10000\ |z|<1\}$ and $\Delta^{\ast}\coloneqq\Delta\setminus\{0\}$. Further, let $f:\mathcal{X}\rightarrow\Delta$ be a semistable degeneration with central fibre $X_{0}$ which is a complete simple normal crossing divisor. Let the generic fibre $X_{t},t\in\Delta^{\ast}$ be smooth and of projective dimension $n$. By Deligne, once $X_{0}$ is a complete simple normal crossing variety, the $H:=H^{*}(X_{0};\mathds{Q})$ can be equipped with a natural mixed Hodge structure coming from the Hodge structures provided by its normalization and by its irreducible components. On the other hand, Schmid constructs a mixed Hodge structure on $X_{0}$ using the information encoded in its smoothing. We now recall the main facts (see [Sch73]).
Let $t_{0}$ be a smooth point, and $T:H^{*}(X_{t_{0}};\mathds{Q})\rightarrow H^{*}(X_{t_{0}};\mathds{Q})$ the Picard-Lefschetz transformation of the family. By Borel’s theorem, $T$ is a quasi-unipotent operator. Let $0<m\in\mathds{Z}$ be the smallest positive integer such that $T^{m}$ is unipotent. We define $N:=\frac{1}{m}\log T^{m}$ as the nilpotent logarithm of monodromy. By Landman’s theorem ([Lan73]), we have $N^{n+1}=0$. Let $\Gamma$ be the monodromy group of the family,
$\phi:\Delta^{*}\rightarrow\Gamma\backslash D\quad\text{and }\widetilde{\phi}:\mathcal{H}\rightarrow D$
be respectively the period map (over $\Delta^{*}$) and a lift to its universal covers. Then the function
$\psi(z):=e^{-zN}F(z),$
satisfies the periodicity conditions
$\psi(z+1)=\psi(z),$
and hence descends to a well-defined function $\psi(s)$ on $\Delta^{*}$ via the covering map $\mathcal{H}\to\Delta^{*},z\mapsto s=e^{2\pi iz}.$ Then, Schmid’s Nilpotent orbit theorem assert that $\psi$ extends holomorphically over $0\in\Delta$, and the nilpotent orbit $\theta(z):=e^{zN}\cdot F_{\infty}$, with $F_{\infty}:=\psi(0)$, asymptotically approximates $\phi$. Moreover, the nilpotent orbit yields a limiting mixed Hodge structure as the data of the Hodge structure
$(H_{\mathds{Q}},F_{\infty},W(N)),$
where $W(N)_{*}$ is the monodromy weight filtration defined as the unique rational increasing filtration such that
$N(W(N)_{l})\subset W(N)_{l-2},\forall,l\quad\text{and}\,N^{a}:\operatorname{Gr}^{W(N)}_{a}\xrightarrow{\cong}\operatorname{Gr}^{W(N)}_{-a},\forall\,a\geq 0.$
In other words, the family $\phi(t)$ of Hodge structure on $H^{n}(X_{t},\mathds{Q})$ degenerates to the mixed Hodge structure $(F_{\infty},W(N))$ as $t\rightarrow 0$. Notice that $F_{\infty}$ depends on the choice of the coordinate, while $N,\theta(z)$ are independent of it.
In the case of a degeneration of mixed Hodge structure, the picture is more complicated since the limit mixed Hodge structure exists if the variation is admissible, i.e. the Hodge filtration $F^{\bullet}_{t}$ extends, and the relative weight filtration $M:=M(N,W)$ exists, where $M$ satisfies
$N(M_{l})\subset M_{l-2},\forall,l\quad\text{and}\,N^{l}:\operatorname{Gr}^{M}_{l+k}\xrightarrow{\cong}\operatorname{Gr}^{M}_{k-l},\forall\,k,l\geq 0.$
Every geometric variation is always admissible, and so, by uniqueness, $W(N)$ and $M$ coincide (up to a shift). Finally, we recall that for such a limit MHS, the operator $N$ becomes a (1,1)(-1,-1)-morphism.
Deligne and Schmid’s results are connected by the Clemens-Schmid exact sequence, which works as follows.
Theorem 2.1.
Let $\mathcal{X}\rightarrow\Delta$ be a semistable degeneration.
  1. (i)
    There exists a retraction $\mathcal{X}\rightarrow X_{0}$ which provides $H^{n}(\mathcal{X},\mathds{Q})\cong H^{n}(X_{0},\mathds{Q})$ with a mixed Hodge structure.
  2. (ii)
    There is the exact sequence of morphisms of mixed Hodge structures
    $\ldots H_{2d+2-n}(\mathcal{X};\mathds{Q})\xrightarrow{\alpha}H^{n}(\mathcal{X};\mathds{Q})\xrightarrow{i^{*}}H^{n}(\mathcal{X}_{t_{0}};\mathds{Q})\xrightarrow{N}H^{n}(\mathcal{X}_{t_{0}};\mathds{Q})\xrightarrow{\beta}H_{2d-n}(\mathcal{X};\mathds{Q})\xrightarrow{\alpha}\ldots.$
As a consequence, cocycles invariant under monodromy, i.e. cocycles in $\ker N$, are cocycles on the special fibre. Furthermore, $N^{n+1}=0$ always and $N^{n}=0$ if and only if $H^{n}(|\Gamma(X_{0})|)=0$, where $|\Gamma(X_{0})|$ is the support of the dual graph of $X_{0}$.
2.2. Degeneration of smooth projective curves and MHS
Given a degeneration as in subsection 1.2.3, we now study the limit mixed Hodge structure $H^{k}_{lim}(JC_{t};\mathds{Q})$. For the sake of completeness, we start from the structure of $H^{1}_{lim}(C_{t};\mathds{Q})$, even if it is well-known in the literature (for example, see [Ca80]).
2.2.1. The case of central irreducible nodal curve
We consider the first degeneration described in subsection 1.2.3: Here, the central fibre over $0\in\Delta$ is an irreducible nodal curve $C_{0}$, the node $x_{0}$ is obtained by identifying two points $p,q$ in the normalization $\widetilde{C}_{0}$. Furthermore, the cycle pqp-q is torsion. The natural variation of Hodge structure, whose fibre over $t\in\Delta$ is provided by $H^{1}(C_{t},\mathbb{Q})$, has a limit MHS, as we now recall. Let $H_{1}(C_{t},\mathbb{Z})=\langle\alpha_{1},\alpha_{2},\alpha_{3},\beta_{1},\beta_{2},\beta_{3}\rangle$ be a symplectic basis (similarly $\langle\alpha^{*}_{1},\alpha^{*}_{2},\alpha^{*}_{3},\beta^{*}_{1},\beta^{*}_{2},\beta^{*}_{3}\rangle$ the dual basis in cohomology), and $\alpha_{3}$ the vanishing cycle of the degeneration. Then, since $N^{2}=0$, the monodromy weight filtration has 3 weights and
$M_{0}(H^{1}_{lim}(C_{t};\mathds{Q}))=\operatorname{Im}N$
$M_{1}(H^{1}_{lim}(C_{t};\mathds{Q}))=\operatorname{ker}N=H^{1}(C_{0};\mathds{Q})$
$M_{2}(H^{1}_{lim}(C_{t};\mathds{Q}))=H^{1}_{lim},$
where the equality in $M_{1}$ follows from Theorem 2.1, since up to a base change, our family admits a semistable degeneration. We want the mixed Hodge structure on $H^{1}_{lim}(C_{t};\mathds{Q})$ to be split, namely we want $H^{1}_{lim}(C_{t};\mathds{Q})=\oplus_{k}\operatorname{Gr}^{M}_{k}H^{1}_{lim}(C_{t};\mathds{Q})$, where
$\operatorname{Gr}^{M}_{0}H^{1}_{lim}(C_{t};\mathds{Q})=\mathds{Q}\beta_{g}^{*},\;V=\operatorname{Gr}^{M}_{1}H^{1}_{lim}(C_{t};\mathds{Q})=H^{1}(\widetilde{C}_{0};\mathds{Q}),\;\text{and}\operatorname{Gr}^{M}_{2}H^{1}_{lim}(C_{t};\mathds{Q})=\mathds{Q}(-1)\alpha_{g}^{*}.$
Indeed, the extension of $\operatorname{Gr}^{M}_{1}$ by $\operatorname{Gr}^{M}_{0}$ belongs to $\operatorname{Ext^{1}}(\mathbb{Q},H^{1}(\widetilde{C}_{0};\mathds{Q}))\cong J\widetilde{C}_{0}\otimes\mathds{Q}$, and it is given by the class of pqp-q. Since by assumption 1.8 pqp-q is torsion, it is zero in $J\widetilde{C}_{0}\otimes\mathds{Q}$. The same (dually) happens to the extensions of $\operatorname{Gr}^{M}_{2}$ by $\operatorname{Gr}^{M}_{1}$. Finally, the extension of $\operatorname{Gr}^{M}_{2}$ by $\operatorname{Gr}^{M}_{0}$ is an element $\xi\in\operatorname{Ext}^{1}(\mathds{Q}(-1),\mathds{Q})\cong\mathds{C}^{*}\otimes\mathds{Q}$. One can choose a coordinate for $\Delta$ to make this extension trivial: Let $s=ft$ be a new choice of coordinate. Then $\log s=\log(f)+\log t$, and $F_{s,\infty}=e^{-\frac{1}{2\pi i}\log(f(0))}F_{t,\infty}$. Let $c=\frac{1}{2\pi i}\log f(0)$. This new choice of coordinate changes $\xi$ to $\xi-c$. Now, we can choose a coordinate such that $c=\xi$, which makes the above extension trivial.
Assumption 2.2.
We assume that the parameter $t\in\Delta$ is such that the extension $\xi\in\operatorname{Ext}^{1}(\mathds{Q}(-1),\mathds{Q})$ is trivial.
With the above assumption, we obtain
$H^{1}_{lim}(C_{t};\mathds{Q})=\mathbb{Q}\oplus H^{1}(\widetilde{C}_{0};\mathbb{Q})\oplus\mathbb{Q}(-1)$
as a split mixed Hodge structure. Now, using the fact that $H^{k}(JC_{t};\mathds{Q})\cong\wedge^{k}H^{1}(C_{t};\mathds{Q})$, we compute $H^{k}_{lim}(JC_{t}:\mathbb{Q})$, by taking the exterior power $\wedge^{k}H^{1}_{lim}(JC_{t};\mathbb{Q})\cong\wedge^{k}H^{1}_{lim}(C_{t};\mathbb{Q})$. We keep calling the nilpotent logarithm of monodromy $\wedge^{k}N$ of the variation $H^{k}(JC_{t};\mathds{Q})$ by $N$, satisfying $N^{2}=0$.
It follows that the limit MHS $H^{k}_{lim}(JC_{t};\mathbb{Q})$ has 3 weights with respect to relative weight filtration $M$, with graded pieces:
$\operatorname{Gr}^{M}_{k-1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0},\mathds{Q}),$
$\operatorname{Gr}^{M}_{k}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-2}(J\widetilde{C}_{0};\mathds{Q}(-1))\oplus H^{k}(J\widetilde{C}_{0};\mathds{Q}))\cong H^{k}(\widetilde{X}_{0};\mathds{Q}),$
$\operatorname{Gr}^{M}_{k+1}H^{k}_{lim}(JC_{t};\mathbb{Q})=H^{k-1}(J\widetilde{C}_{0};\mathds{Q}(-1)).$
The isomorphism in the middle follows from the projective bundle formula, noting that $\widetilde{X}_{0}$ is $\mathbb{P}^{1}$-bundle over $J\widetilde{C}_{0}$.
2.2.2. The case of central reducible curve
We now consider the second degeneration in 1.2.3: Here the central fibre over $0\in\Delta$ is a reducible curve $C_{0}=C_{1}\cup C_{2}$ obtained as the union of two irreducible smooth components $C_{1}$ and $C_{2}$ of genera $1$ and $2$ respectively, with a separating node ${y_{0}}=C_{1}\cap C_{2}$. As before, we first consider the VHS provided by $H^{1}(C_{t};\mathds{Q})$, for $t\in\Delta^{*}$. In this case, $N=0$ and the limit MHS trivially is as follows:
$M_{0}(H^{1}_{lim}(C_{t};\mathds{Q}))=\operatorname{Im}N=\{e\}$
$M_{1}(H^{1}_{lim}(C_{t};\mathds{Q}))=\operatorname{ker}N=H^{1}(C_{0};\mathds{Q})=H^{1}(C_{1};\mathds{Q})\oplus H^{1}(C_{2};\mathds{Q}).$
Therefore, the equality
$H^{k}_{\lim}(JC_{t};\mathbb{Q})$ $=\oplus_{r+s=k}\wedge^{r}H^{1}(C_{1};\mathds{Q})\otimes\wedge^{s}H^{1}(C_{2};\mathds{Q})$
$=H^{k}(JC_{2};\mathds{Q})\oplus H^{1}(JC_{1};\mathds{Q})\otimes H^{k-1}(JC_{2};\mathds{Q})\oplus H^{k-2}(JC_{2};\mathds{Q}(-1))$
$\cong H^{k}(J(C_{1})\times J(C_{2});\mathds{Q}),$
provides $H^{k}_{\lim}(JC_{t};\mathbb{Q})$ with a pure Hodge structure of weight $k$.
Remark 2.3.
Since $N=0$, this is the case for no degeneration, and we technically do not need assumption 2.2. However, we work with it to bring both the degenerations under a common set up.
3. Biextension mixed Hodge structure and its variation
Here, we introduce the mixed Hodge structure that is the main focus of the article.
Definition 3.1.
Let $H$ be a rational pure Hodge structure of weight 1-1. A biextension $B_{H}$ associated with $H$ is a mixed Hodge structure that has at most three non trivial graded pieces, satisfying
$\operatorname{Gr}_{k}^{W}(B_{H})=\begin{cases}\mathds{Q}(0),&\text{ if }k=0,\\ H,&\text{ if }k=-1,\\ \mathds{Q}(1),&\text{ if }k=-2,\\ 0,&\text{ otherwise.}\end{cases}$
The set $\mathcal{B}_{H}$ of all biextensions associated with $H$ is a $\mathds{C}^{\ast}\otimes\mathds{Q}$ bundle over $J(H)\times J(H)^{\vee}$, where $J(H)=\operatorname{Ext}^{1}(\mathds{Q}(0),H)$ is the rational Jacobian associated to $H$ and $J(H)^{\vee}=J(H^{\vee})=\operatorname{Ext}^{1}(H^{\vee},\mathds{Q}(1))$ is its dual. Since $J(H)\otimes\mathds{R}=J(H)^{\vee}\otimes\mathds{R}=0$, one can identify
$\mathcal{B}_{H}\otimes\mathds{R}\cong\operatorname{Ext}^{1}(\mathds{R}(0),\mathds{R}(1))\cong\mathds{R}.$
Hence, to any biextension $B_{H}$, one can attach a unique real number $\operatorname{ht}(B_{H})$ called the (archimedean) height of $B_{H}$. Notice that (cf. Remark 1.6) for a biextension mixed Hodge structure $\delta=\delta_{-2}=\delta^{-1,-1}$ with respect to $\text{ad}(Y)$, since there are no eigencomponents of degree 0 or 1-1. Then it is easy to see that
Theorem 3.2.
Given a biextension $B_{H}$ associated to a pure Hodge structure $H$ of weight 1-1, let $1_{B_{H}}\in\operatorname{Gr}^{W}_{0}B_{H}=\mathds{Q}(0)$ and $1^{\vee}_{B_{H}}\in\operatorname{Gr}^{W}_{-2}B_{H}=\mathds{Q}(1)$ be the canonical (Betti) generators. Then
$\delta_{-2}(1_{B_{H}})=\frac{1}{2\pi}\operatorname{ht}(B_{H})1^{\vee}_{B_{H}}.$
Remark 3.3.
From the above theorem, it is immediate that $\operatorname{ht}(B_{H})=0$ if and only if $B_{H}$ is split over $\mathds{R}$, that is $\overline{I}^{a,b}_{B_{H}}=I^{b,a}_{B_{H}}$ for all pairs $(a,b)\in\mathds{Z}\times\mathds{Z}$. In this sense, the height is an obstruction to $\mathds{R}$-splitness of a biextension.
Next, we restate the relation between heights and a particular type of morphisms of mixed Hodge structures.
Proposition 3.4.
Let $B_{1}$ and $B_{2}$ be biextension mixed Hodge structures, with $1_{B_{1}}\in Gr^{W_{1}}_{0}B_{1}=\mathds{Q}(0)$ and $1^{\vee}_{B_{1}}\in\operatorname{Gr}^{W_{1}}_{-2}=\mathds{Q}(1)$ be the canonical (Betti) generators (respectively $1_{B_{2}}$ and $1^{\vee}_{B_{2}}$ for $B_{2}$). Let $f\colon(B_{1},F_{1},W_{1})\to(B_{2},F_{2},W_{2})$ be a morphism of mixed Hodge structures such that $f(1_{B_{1}})=1_{B_{2}}$ and $f(1^{\vee}_{B_{1}})=1^{\vee}_{B_{2}}$. Then $\operatorname{ht}(B_{1})=\operatorname{ht}(B_{2})$.
Proof.
The proof is exactly similar to [BGP22, Proposition 2.8] in a relatively easier setting. We restate it for the sake of completeness. Using Lemma 1.7, and the fact that $\delta_{B_{i}}=\delta_{B_{i},-2}$ for $i=1,2$, we have $f\circ\delta_{B_{1},-2}=\delta_{B_{2},-2}\circ f$. Hence
$\frac{1}{2\pi}\operatorname{ht}(B_{2})1^{\vee}_{B_{2}}=\delta_{B_{2},-2}(1_{B_{2}})$ $=\delta_{B_{2},-2}(f(1_{B_{1}}))$
$=f(\delta_{B_{1},-2}(1_{B_{1}}))$
$=\frac{1}{2\pi}\operatorname{ht}(B_{1})f(1^{\vee}_{B_{1}})$
$=\frac{1}{2\pi}\operatorname{ht}(B_{1})1^{\vee}_{B_{2}}.$
Now it’s immediate that $\operatorname{ht}(B_{2})=\operatorname{ht}(B_{1})$. ∎
3.1. Geometric biextension
In geometric settings, biextension mixed Hodge structures arise as relative cohomology groups as follows. Let $X$ be a smooth, projective, and complex variety of dimension $d$, $Z$ and $W$ two algebraic cycles of codimensions $p$ and $q$ respectively that are homologous to zero and intersect properly. Let us also assume the numerical relation between the codimensions $p+q=d+1$ (together with the proper intersection, which means that $|Z|\cap|W|=\emptyset$). For ease of exposition, we also assume that $Z=Z_{1}-Z_{0}$ and $W=W_{1}-W_{0}$ for reduced and irreducible subvarieties $\{Z_{0},Z_{1},W_{0},W_{1}\}$. Let $H=H^{2p-1}(X;\mathds{Q}(p))$, which is a pure Hodge structure of weight 1-1. Using the long exact sequences associated with cohomology with support and relative cohomology, together with purity, and the fact that $|Z|\cap|W|=\emptyset$, we get the commutative diagram 1.
[…]
Figure 1. Biextension diagram
In the diagram, we have
$\mathds{Q}(0)=\ker\left(H^{2p}_{|Z|}(X;\mathds{Q}(p))\to H^{2p}(X;\mathds{Q}(p))\right),$
and
$\mathds{Q}(1)=\text{Coker}\left(H^{2p-2}(X;\mathds{Q}(p))\to H^{2p-2}(|W|;\mathds{Q}(p))\right),$
noticing the relation $2p-2=2(d+1-q)-2=2\dim(W)$. The top horizontal exact sequence in diagram 1 gives the extension class for $Z$ and equals the Abel-Jacobi image of $Z$, while the left vertical exact sequence gives the dual of the extension class for $W$ twisted by $\mathds{Q}(1)$.
From the functoriality of mixed Hodge structures, it is easy to see that $B_{Z,W}\coloneqq H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))$ satisfies the condition on the graded pieces to be a biextension. We remark that in case the cycles $Z$ and $W$ have more components, then one gets a subquotient of the mixed Hodge structure $H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))$ as a biextension.
Remark 3.5.
Using the empty intersection between $|Z|$ and $|W|$ once again, we get the duality
$H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))\cong H^{2q-1}(X\setminus|W|,|Z|;\mathds{Q}(q))^{\vee}.$
This can be used to show that $B^{\vee}_{Z,W}\cong B_{W,Z}$, where $B_{W,Z}$ is the biextension obtained with the roles of $Z$ and $W$ interchanged.
Remark 3.6.
In case of a biextension of geometric origin as above, the corresponding height $\operatorname{ht}(B_{Z,W})$ equals the archimedean height pairing $\langle Z,W\rangle$ between the two cycles (this is a consequence of [Ha90, Proposition 3.2.13]. See also [Pe06, §5] for further insights.). In fact, using the duality of remark 3.5 one can interpret $\operatorname{ht}(B_{Z,W})$ as $\int_{Z}\mathfrak{g}_{W}$, where ${\mathfrak{g}}_{W}$ is a Green form of logarithmic type for the cycle $W$ that can be shown to exist. For example, in $\mathds{P}^{1}$ any choice of four distinct points $\{p,q,r,s\}$ gives rise to homologically (actually rationally) trivial cycles Z=pqZ=p-q and W=rsW=r-s. In this case, one choice for ${\mathfrak{g}}_{W}$ is the function $\log\Big|\frac{t-r}{t-s}\Big|^{2}$ that has logarithmic singularities on $|W|=\{r,s\}$. The height is given by the logarithm of the cross-ratio of the four points.
3.2. Abstract degeneration of biextension variations
Our main reference for this is [BP19, §2, §3]. We point out that the treatment in [BP19] holds more generally for an admissible nilpotent orbit with values in the classifying space of graded-polarized mixed Hodge structures. Here we restrict to admissible, graded-polarized biextension variations over $\Delta^{\ast}$.
Definition 3.7.
Let $\mathcal{H}$ be a weight 1-1 polarized variation of rational pure Hodge structures over $\Delta^{\ast}$ with unipotent monodromy. We define a biextension variation $\mathcal{B}_{\mathcal{H}}$ associated with $\mathcal{H}$ as an admissible variation of mixed Hodge structures over $\Delta^{\ast}$, with isomorphisms
$\operatorname{Gr}_{k}^{W}(\mathcal{B}_{\mathcal{H}})=\begin{cases}\mathds{Q}(0),&\text{ if }k=0,\\ \mathcal{H},&\text{ if }k=-1,\\ \mathds{Q}(1),&\text{ if }k=-2,\\ 0,&\text{ otherwise.}\end{cases}$
For each $t\in\Delta^{\ast}$, one can associate the height $\operatorname{ht}(\mathcal{B}_{\mathcal{H},t})$, which defines a $C^{\infty}$-function
$h\colon\Delta^{\ast}\to\mathds{R},\penalty 10000\ h(t)\coloneqq\operatorname{ht}(\mathcal{B}_{\mathcal{H},t}).$
This function does not always extend continuously to $\Delta$. There is a $\mu\in\mathds{Q}$, such that the amended function
$\widetilde{h}(t)=h(t)+\mu\log|t|$
has a continuous extension. For the rest of this section, we briefly discuss how one obtains this $\mu$ and a limit height.
Consider the biextension variation $\mathcal{B}_{\mathcal{H},t}=(F(t),W),\penalty 10000\ t\in\Delta^{\ast}$. As in section 2.1, let $N$ be the monodromy operator, $\mathcal{B}_{\mathcal{H},\lim}=(F_{\infty},M)$ the limit mixed Hodge structure with relative weight filtration $M=M(N,W)$, and $(N,F_{\infty},W)$ the admissible nilpotent orbit. Also, let $\delta_{M}$ be the Deligne splitting associated to the limit mixed Hodge structure $(F_{\infty},M)$, and $\widetilde{F}_{\infty}=e^{-i\delta_{M}}F_{\infty}$ be the split-Hodge filtration. Then, there exists a unique functorial $\mathds{R}$-grading of the original weight filtration $W$ defined first by Deligne (see [Pe01, Theorem 4.4], [KP03, §3,§4], and [BP19, Lemma 65] for details)
$Y(N,F_{\infty},W)=\text{Ad}(e^{-iN})Y_{(e^{iN}\widetilde{F}_{\infty},W)},$
which commutes with $Y_{(\widetilde{F}_{\infty},M)}$.
Remark 3.8.
For admissible graded-polarized biextension variations, one has ([Pe06, Theorem 4.15])
$Y(N,F_{\infty},W)=\lim_{\operatorname{Im}(z)\to\infty}e^{-zN}\cdot Y_{(F(z),W)}.$
Since $Y(N,F_{\infty},W)$ is a grading of the original weight filtration, one has the decompositions of $N$ and $\delta_{M}$ into eigencomponents modulo $\text{ad}(Y(N,F_{\infty},W))$, with respect to the filtration $W$:
$N=N_{0}+N_{-1}+N_{-2},$
$\delta_{M}=\delta_{M,0}+\delta_{M,-1}+\delta_{M,-2},$
where $\text{ad}(Y(N,F_{\infty},W))(N_{-j})=-jN_{-j}$, resp. $\text{ad}(Y(N,F_{\infty},W))(\delta_{M,-j})=-j\delta_{M,-j}$, for $j=0,1,2$ (since these are the weights for $W$). In fact, since we are working with a biextension variation, we have
$\operatorname{Gr}_{k}^{W}(\mathcal{B}_{\mathcal{H},\lim})=\begin{cases}\mathds{Q}(0),&\text{ if }k=0,\\ \mathcal{H}_{\lim},&\text{ if }k=-1,\\ \mathds{Q}(1),&\text{ if }k=-2,\\ 0,&\text{ otherwise},\end{cases}$
with respect to the original weight filtration $W$. Hence the components $N_{-2}$ are $\delta_{M,-2}$ are easy to extract: $N_{-2}$ is the component of $N$ that goes from $\mathds{Q}(0)$ to $\mathds{Q}(1)$, while $\delta_{M,-2}$ is the component of $\delta_{M}$ that goes from $\mathds{R}(0)$ to $\mathds{R}(1)$.
Remark 3.9.
Notice that, $\mathcal{B}_{\mathcal{H},\lim}$ may not have a biextension structure with respect to the relative weight filtration $M$. Also, the components $\delta_{M,0}$ and $\delta_{M,-1}$ can be more complex, depending on the degree of vanishing of the monodromy operator $N$.
Let $1$ denote the canonical Betti generator of $\mathds{Q}(0)$ and $1^{\vee}$ that of $\mathds{Q}(1)$. Following [Pe06, Theorem 5.19], we define $\mu$ by
(3.1) $N_{-2}(1)=\mu 1^{\vee}.$
Since $N$ does not depend on the choice of a coordinate for $\Delta$, neither does $\mu$. On the other hand, the limit height is defined by ([BP19, (71)])
(3.2) $\delta_{M,-2}(1)=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee}.$
In this case, since the limit mixed Hodge structure depends on a choice of coordinate for $\Delta$, so does the height $H(N,F_{\infty},W)$. The following result justifies the name ‘limit height’ ([BP19, Theorem 76]).
Proposition 3.10.
The asymptotic height $\tilde{h}(t)=h(t)+\mu\log|t|$ extends continuously to $t=0$, with $\tilde{h}(0)=H(N,F_{\infty},W)$.
4. Degeneration of a family of Ceresa cycles and their heights
In this section, given a family of genus three curves $\mathcal{C}\to\Delta$, we introduce (a family of) Ceresa cycles, and describe their behaviour in the limit.
4.1. Degeneration of Ceresa cycles
Let $\mathcal{C}\rightarrow\Delta$ be a family as in 1.2.3. We first discuss the case when $C_{0}$ is an irreducible nodal curve, with one node $x_{0}$.
Let $\beta:\Delta^{*}\to\mathcal{C},t\to b_{t}$ be a section fixing a base point $b_{t}\in C_{t},t\neq 0$, and consider a holomorphic extension to $\Delta$, sending $t=0$ to a point $b\in C_{0},b\neq x_{0}$. For $t\in\Delta^{*}$, the Abel Jacobi map (1.2.2) embeds $C_{t}\hookrightarrow JC_{t}$ through $x\mapsto x-b_{t}$, and the Ceresa cycle is defined by
$Z_{t}:=C_{t}-[-1]^{\ast}C_{t}\in Z^{2}(JC_{t}).$
In general, this cycle gives a non-trivial element in the codimension two Griffiths group of $JC_{t}$. Using results in section 1.2, we get what follows.
Proposition 4.1.
The family of Ceresa cycle $Z_{t}\in Z^{2}(JC_{t})$ extends to $Z_{0}\coloneqq C_{0}-[-1]^{\ast}C_{0}$ in $Z^{2}(X_{0})$, which is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$.
Proof.
The first statement follows from (1.7), which embeds $C_{0}$ in $X_{0}$, and from (1.4), which implies that the [1][-1] morphism can be lifted to $X_{0}$. Next, using standard Verdier duality isomorphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong H^{0}(Z_{0};\mathds{Q}(0))$ and the fact that $|Z_{0}\cap S|=\{(p-b,0),(b-p,\infty)\}$ is finite, we get that $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\cong\mathds{Q}(0)\oplus\mathds{Q}(0)$. The morphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))$ defines the cycle class map. We show that $Z_{0}$ is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$. The cycle class of $Z_{0}$ lands in the weight zero component of $H^{4}(X_{0};\mathds{Q}(2))$ which by equation (1.6) is given by $H^{4}(\widetilde{X}_{0};\mathds{Q}(2))$. Further, since $Z_{0}|_{JC_{0}}=0$ in $\operatorname{CH}^{2}(JC_{0})$, we use the localization exact sequence of Chow groups
$\operatorname{CH}^{1}(J\widetilde{C}_{0})\xrightarrow{\iota_{*}}\operatorname{CH}^{2}(X_{0})\rightarrow\operatorname{CH}^{2}(JC_{0}),$
to conclude that $Z_{0}=\iota_{*}(\Gamma)$, for $\Gamma=[p-q]-[q-p]\in\operatorname{CH}^{1}(J(\widetilde{C}_{0}))$. Hence, the cycle class of $Z_{0}$ belongs to $H^{4}(X_{0};\mathds{Q}(2))^{-}$, which is the anti-invariant subspace of $H^{4}(X_{0};\mathds{Q}(2))$ with respect to the involution morphism [1][-1]. But the involution [1][-1] acts as identity on $H^{4}(\widetilde{X}_{0};\mathds{Q}(2))$, implying $H^{4}(\widetilde{X}_{0};\mathds{Q}(2))^{-}=0$. Hence, we can conclude the result. ∎
Remark 4.2.
In the normalization $\widetilde{X}_{0}$ of $X_{0}$, the class of the cycle $\widetilde{Z}_{0}$ (the normalization of $Z_{0}$) has an explicit description using the projective bundle formula. Indeed, recall that $\widetilde{X}_{0}=\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L_{p-q})$ is the projective bundle over $J\widetilde{C}_{0}$. Under the projective bundle isomorphism
$\operatorname{CH}^{2}(\widetilde{X}_{0})\cong\operatorname{CH}^{2}(J\widetilde{C}_{0})\oplus\operatorname{CH}^{1}(J\widetilde{C}_{0}),$
the image of the cycle $\widetilde{Z}_{0}$ is given by $(p-q,\widetilde{C}_{0}-[-1]^{\ast}\widetilde{C}_{0})$, where we view pqp-q as a zero cycle in $J\widetilde{C}_{0}$. Since by assumption 1.8 the zero cycle pqp-q is torsion, the class of $\widetilde{Z}_{0}$ is trivial in $\operatorname{CH}^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. Hence $\widetilde{Z}_{0}$ is rationally equivalent to zero in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$.
Even though $X_{0}$ is a singular variety, we show in the next proposition that purity holds for the pair $(X_{0},Z_{0})$. This is probably well known to experts, but we provide a proof for completeness.
Proposition 4.3.
The cohomology with support for the pair $(X_{0},Z_{0})$ satisfies
$H^{k}_{|Z_{0}|}(X_{0};\mathds{Q}(2))=\begin{cases}0,&\text{ if }k<4,\\ \mathds{Q}(0)\oplus\mathds{Q}(0)&\text{ if }k=4.\end{cases}$
Moreover, $\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)\cong\mathds{Q}(0)$, generated by the fundamental class $\{Z_{0}\}$.
Proof.
The equality for $k=4$ is already shown in Proposition 4.1. For ease of writing, we rename $U=JC_{0}$, which is the open component of $X_{0}$, $S=J\widetilde{C}_{0}$, the singular locus, and $Z_{0,U}\coloneqq Z_{0}\cap U$. We will use the following standard facts: From definition, the cohomology with compact support is given by $H^{k}_{c}(U;\mathds{Q}(2))=H^{k}(X_{0},S;\mathds{Q}(2))$, and $H^{k}_{|Z_{0,U}|,c}(U;\mathds{Q}(2))=H^{k}_{|Z_{0,U}|}(U;\mathds{Q}(2))$. Now, we specialize the long exact sequence of relative cohomology with support in $|Z_{0}|$
$\cdots\to H^{k-1}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k}_{|Z_{0,U|}}(U;\mathds{Q}(2))\to H^{k}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\\ \to H^{k}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))\to H^{k+1}_{|Z_{0,U}|}(U;\mathds{Q}(2))\to\cdots,$
to $k<4$. Since $S$ is smooth and projective and $|Z_{0}\cap S|$ is finite, we can use semi-purity to conclude that $H^{k-1}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))=H^{k}_{|Z_{0}\cap S|}(S;\mathds{Q}(2))=0$. Since $U$ is smooth, semi-purity once again asserts that $H^{k}_{|Z_{0,U}|}(U;\mathds{Q}(2))=0$. Hence, we conclude $H^{k}_{|Z_{0}|}(X_{0};\mathds{Q}(2))=0$ for $k<4$. For the last statement, we use the fact that $Z_{0}$ is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$. This implies that $\operatorname{cl}(C_{0})=\operatorname{cl}([-1]^{\ast}C_{0})$ in $H^{4}(X_{0};\mathds{Q}(2))$. Notice that, every element in $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))$ is of the form $\xi=x\{C_{0}\}+y\{[-1]^{\ast}C_{0}\}$ for $x,y\in\mathds{Q}$, where $\{Y\}$ denotes the fundamental class of a cycle $Y$ in $H^{k}_{|Y|}(X;\mathds{Q}(p))$. The morphism $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q})$ is defined by the assignment $\{C_{0}\}\to\operatorname{cl}(C_{0})$ (respectively $\{[-1]^{\ast}C_{0}\}\to\operatorname{cl}([-1]^{\ast}C_{0})$). Hence we can conclude using $\operatorname{cl}(C_{0})=\operatorname{cl}([-1]^{\ast}C_{0})$ that $x+y=0$, and
$\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)=\langle\{Z_{0}\}\rangle\cong\mathds{Q}(0).$
Let us move to the second degeneration discussed in section 1.2.3: We have a family $\mathcal{C}\rightarrow\Delta$ of genus three curves, whose central fibre $C_{0}=C_{1}\cup C_{2}$ is reducible, with smooth irreducible components $C_{1}$ of genus $1$ and $C_{2}$ of genus $2$, meeting at $y_{0}=C_{0}\cap C_{1}$. In this case, the following holds:
Proposition 4.4.
The family of Ceresa cycles $Z_{t}\in Z^{2}(JC_{t})$ degenerates to a cycle ${Z}_{0}:=C_{0}-C_{0}^{-}\in JC_{0}$ in $Z^{2}(JC_{0})=Z^{2}(JC_{1}\times JC_{2})$ which is homologous to zero.
Proof.
The image of $C_{0}\subset JC_{0}$ via the Abel Jacobi map (1.2.3) is given by two irreducible components meeting at the image of the node. Let us call the two components $C_{1},C_{2}$. Furthermore, let us observe that, since $JC_{0}\cong JC_{1}\times JC_{2}$, the morphism [1][-1] of $JC_{0}$ corresponds to the multiplication ([1],[1])([-1],[-1]) on $JC_{1}\times JC_{2}$. Let $q_{i}\in C_{i},i=1,2$ be the preimages of the node $y_{0}\in C_{0}$, and, to simplify the argument, let us assume that $x\in C_{2}$ is a smooth Weiestrass point. Hence,
$Z_{0}$ $=\mathcal{O}_{E}\times(C_{2}-[-1]^{*}C_{2})+C_{1}\times\mathcal{O}_{C_{2}}(q_{2}-x)-[-1]^{*}C_{1}\times\mathcal{O}_{C_{2}}(x-q_{2})$
Since $JC_{2}$ is an abelian surface, $C_{1}\times\mathcal{O}_{C_{2}}(q_{2}-x)$ and $C_{1}\times\mathcal{O}_{C_{2}}(x-q_{2})$ are homotopic, using as always that [1][-1] acts as the identity on even homology, we can conclude the result. A similar argument applies when $x$ is not a Weiestrass point. ∎
4.2. Degeneration of height pairings
We now consider the geometric biextension recalled in section 3. We study this construction for a family $\mathcal{C}\to\Delta$ of genus three curves, and analyze the limits of the associated heights. We consider the two degenerations of the Ceresa cycles described in subsection 4.1.
Let us first discuss the case when the fibre $C_{0}$ over $0\in\Delta$ is nodal and irreducible. Let us consider two sections $\beta_{i}:\Delta\to\mathcal{C}$, $i=1,2$, with $b_{1,t}:=\beta_{1}(t)\neq\beta_{2}(t)=:b_{2,t}$, and for $t\neq 0$ let $C_{t,b_{i,t}}$ be the image of the Abel Jacobi embeddings. Let
(4.1) $Z_{t}:=C_{t,b_{1,t}}-[-1]^{*}C_{t,b_{1,t}},\quad W_{t}:=C_{t,b_{2,t}}-[-1]^{*}C_{t,b_{2,t}},t\neq 0$
the corresponding Ceresa cycles. Set $b_{1,0}:=\beta_{1}(0)$ and $b_{2,0}:=\beta_{2}(0)$, and assume that $b_{1,0}\neq b_{2,0}$. Using these points, define
$Z_{0}:=C_{0,b_{1,0}}-[-1]^{*}C_{0,b_{1,0}},\qquad W_{0}:=C_{0,b_{2,0}}-[-1]^{*}C_{0,b_{2,0}},$
as the extensions in $0\in\Delta$ of the families $Z_{t},W_{t}$ as in Proposition 4.1. Then the following holds:
Proposition 4.5.
The Ceresa cycles $Z_{t}$ and $W_{t}$ intersect properly in $JC_{t}$ for $t\in\Delta^{*}$, in other words, they are disjoint. Moreover, the same conclusion holds for $Z_{0}$ and $W_{0}$ in $X_{0}$.
Proof.
Let $t\in\Delta^{*}$. We claim that if
(4.2) $b_{2,t}\notin\{y-x+b_{1,t},x,y\in C_{t}\}\cup\{y+x-b_{1,t},x,y\in C_{t}\},$
then $Z_{t}\cap W_{t}=\emptyset$. Indeed, assume by contradiction that we have $Z_{t}\cap W_{t}\neq\emptyset$. Then either $C_{t,b_{1,t}}\cap C_{t,b_{2,t}}\neq\emptyset$ or $C_{t,b_{1,t}}\cap C_{t,b_{2,t}}^{-}\neq\emptyset$. In the first case $b_{2,t}$ belongs to $\{y-x+b_{1,t},x,y\in C_{t}\}$, while in the second to $\{y+x-b_{1,t},x,y\in C_{t}\}$. The other intersections $C_{t,b_{1,t}}^{-}\cap C_{t,b_{2,t}}^{-},C_{t,b_{1,t}}^{-}\cap C_{t,b_{2,t}}$ give equivalent condition. Therefore, since the locus in (4.2) is of dimension two, and $JC_{t}$ is three-dimensional, there exists $b_{2,t}\in C_{t}$ satisfying (4.2), and thus $Z_{t}\cap W_{t}=\emptyset$. To guarantee that the same holds at the limit, let us recall that $Z_{0}$, resp. $W_{0}$, touches the singular locus of $X_{0}$ only at the image of the node $x_{0}$, and at $1]x_{0}$. Using (1.4), these are $(p-b_{1,0},0)$ amd $(b_{1,0}-p,\infty)$ for $Z_{0}$, and $(p-b_{2,0},0)$ and $(b_{2,0}-p,\infty)$ for $W_{0}$. These are clearly different points, as soon as $b_{1,0}\neq b_{2,0}$. On the other hand, on $C_{0}\setminus\{x_{0}\}$, the Abel-Jacobi map behaves much like the smooth scenario. Therefore, we can use the same argument as before to conclude. ∎
Let us now consider the case when the central fiber of our family $\mathcal{C}\to\Delta$ is a reducible curve $C_{0}=C_{1}\cup C_{2}$. Let $\beta_{i}:\Delta^{*}\to\mathcal{C},$ $i=1,2$, define the base points for the Abel Jacobi embeddings for the Ceresa Cycles $Z_{t},W_{t},t\in\Delta^{*}$ as in (4.1). Consider two holomorphic extensions of the $\beta_{i}$’s to $0\in\Delta$ such that $x:=\beta_{1}(0)\in C_{1}$ and $y:=\beta_{2}(0)\in C_{2}$. Then, let $C_{0,x}$ and $C_{0,y}$ in $JC_{1}\times JC_{2}$ be the two Abel Jacobi embedding of $C_{0}$ obtained using (1.2.3) and (1.2.3). Define
$Z_{0}:=C_{0,x}-[-1]^{*}C_{0,x},\quad W_{0}:=C_{0,y}-[-1]^{*}C_{0,y}$
as the extension of the family of the Ceresa cycle as in Proposition 4.4. As before, the following holds:
Proposition 4.6.
The Ceresa cycles $Z_{t}$ and $W_{t}$ intersect properly in $JC_{t}$, and the same conclusion holds for $Z_{0}$ and $W_{0}$ in $JC_{0}\cong JC_{1}\times JC_{2}$.
Proof.
For $t\neq 0$, the proof is exactly similar to that of Proposition 4.5, and for $t=0$ it follows from the choice of the base points for the Abel Jacobi embeddings. ∎
To state our main result, let us first recall that, in our first degeneration, pulling the cycles $Z_{0},W_{0}$ in $X_{0}$ back to the normalization $\widetilde{X}_{0}$, we get two new cycles $\widetilde{Z}_{0}$ and $\widetilde{W}_{0}$ that still intersect properly, and are rationally equivalent to zero in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. Since $\widetilde{X}_{0}$ is smooth and projective of dimension three, we get another biextension mixed Hodge structure
(4.3) $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}=H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}_{0}|,|\widetilde{W}_{0}|;\mathds{Q}(2)),$
with associated height $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$. In the case of our second degeneration, $X_{0}$ is smooth, and no normalization is necessary. The rest of the section is devoted to the following theorem:
Theorem 4.7.
Let $\mathcal{C}\to\Delta$ be a family of genus three curves with central fibre $C_{0}$ of the two types discussed above. Let $\mathcal{Z}=\{Z_{t}\}_{t\in\Delta^{\ast}}$ and $\mathcal{W}=\{W_{t}\}_{t\in\Delta^{\ast}}$ be two families of Ceresa cycles obtained by choosing the base points $b_{i,t},i=1,2$ as above. Then, for a fixed choice of coordinate (assumption 2.2), the limit height $H(N,F_{\infty},W)$ associated with the biextension variation $B_{Z_{t},W_{t}}$ is given by the Deligne splitting of $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}$. Consequently, we have $H(N,F_{\infty},W)=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$.
Proof.
The key point is to construct (under the assumptions 1.8 and 2.2) the limit mixed Hodge structure associated with the variation $\mathcal{V}_{t}\coloneqq H^{3}(JC_{t}\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ of biextension mixed Hodge structures. Even though conceptually it follows from the functorial properties of limit mixed Hodge structures, we provide the details for clarity.
We start with the case of the irreducible nodal central fiber $C_{0}$. First, let us recall the weight graded pieces of the limit mixed Hodge structure $H^{k}_{lim}(JC_{t};\mathds{Q})$ for $k=3$, twisting everything by $\mathds{Q}(2)$ (to bring the weights in the range [2,0][-2,0]):
$\operatorname{Gr}_{k}^{M}H^{3}_{lim}(JC_{t};\mathbb{Q}(2))=\begin{cases}H^{2}(J\widetilde{C}_{0};\mathds{Q}(1)),&\text{ if }k=0,\\ H^{3}(\widetilde{X}_{0},\mathds{Q}(2)),&\text{ if }k=-1,\\ H^{2}(J\widetilde{C}_{0},\mathds{Q}(2))&\text{ if }k=-2,\\ 0,&\text{ otherwise}.\end{cases}$
To obtain the limit of the biextension variation $B_{Z_{t},W_{t}}$, we need to understand the limit of the different cohomology groups appearing in diagram 1. We first need to compute the limit of the Abel-Jacobi image of $Z_{t}$, resp. $W_{t}$. Let us recall that the Abel-Jacobi image of $Z_{t}$ is given by the (twisted by $\mathds{Q}(2)$) extension class
(4.4) $0\rightarrow H^{3}(JC_{t},\mathbb{Q}(2))\rightarrow H^{3}(JC_{t}\setminus|Z_{t}|;\mathds{Q}(2))\rightarrow\mathds{Q}(0)\rightarrow 0.$
Since the limit mixed Hodge structure is functorial with respect to morphisms of admissible variations of mixed Hodge structures (see [SZ85, Lemma 5.26] for the functoriality with respect to the limit Hodge filtration, while the one for monodromy weight filtration is a consequence of its existence), we obtain
(4.5) $0\rightarrow H^{3}_{lim}(JC_{t},\mathbb{Q}(2))\rightarrow H^{3}_{lim}(JC_{t}\setminus|Z_{t}|;\mathds{Q}(2))\rightarrow\mathds{Q}(0)\rightarrow 0,$
which yields
$\operatorname{Gr}_{k}^{M}H^{3}_{lim}(JC_{t}\setminus|Z_{t}|;\mathds{Q}(2))=\begin{cases}\mathds{Q}(0)\oplus H^{2}(J\widetilde{C}_{0};\mathds{Q}(1)),&\text{ if }k=0,\\ H^{3}(\widetilde{X}_{0};\mathds{Q}(2)),&\text{ if }k=-1,\\ H^{2}(J\widetilde{C}_{0};\mathds{Q}(2)),&\text{ if }k=-2,\\ 0,&\text{ otherwise}.\end{cases}$
We now focus on the $\mathds{Q}(0)$ term above, and we give an alternative interpretation. First, let us observe that
$\mathds{Q}(0)$ $=\ker\Big(H^{4}_{|Z_{t}|}(JC_{t};\mathds{Q}(2))\to H^{4}(JC_{t};\mathds{Q}(2))\Big)_{\lim}$
$\cong\ker\Big(H^{4}_{|Z_{t}|}(JC_{t};\mathds{Q}(2))_{\lim}\to H^{4}_{\lim}(JC_{t};\mathds{Q}(2))\Big).$
From Theorem 2.1, one has morphisms $H^{4}(X_{0};\mathds{Q}(2))\to H^{4}_{\lim}(JC_{t};\mathds{Q}(2))$, resp. $H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}_{|Z_{t}|}(JC_{t};\mathds{Q}(2))_{\lim}$, and a commutative diagram
${H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}(X_{0};\mathds{Q}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}_{|Z_{t}|}(JC_{t};\mathds{Q}(2))_{\lim}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}_{\lim}(JC_{t};\mathds{Q}(2)).}$
By Proposition 4.3, we obtain
$\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)\cong\ker\Big(H^{4}_{|Z_{t}|}(J(C_{t});\mathds{Q}(2))_{\lim}\to H^{4}_{\lim}(J(C_{t});\mathds{Q}(2))\Big).$
In fact, using the normalization morphism $\nu\colon\widetilde{X}_{0}\to X_{0}$ we have the commutative diagram
${H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\{Z_{0}\}\mapsto 0}$${H^{4}(X_{0};\mathds{Q}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{\ast}}$${H^{4}_{|\widetilde{Z}_{0}|}(\widetilde{X}_{0};\mathds{Q}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\{\widetilde{Z}_{0}\}\mapsto 0}$${H^{4}(\widetilde{X}_{0};\mathds{Q}(2)).}$
Hence, we can identify
(4.6) $\mathds{Q}(0)$ $=\ker\Big(H^{4}_{|Z_{t}|}(J(C_{t});\mathds{Q}(2))\to H^{4}(J(C_{t});\mathds{Q}(2))\Big)_{\lim}$
$\cong\ker\Big(H^{4}_{|Z_{t}|}(J(C_{t});\mathds{Q}(2))_{\lim}\to H^{4}_{\lim}(J(C_{t});\mathds{Q}(2))\Big)$
$\cong\ker\Big(H^{4}_{|Z_{0}|}(X_{0};\mathds{Q}(2))\to H^{4}(X_{0};\mathds{Q}(2))\Big)=\langle\{Z_{0}\}\rangle$
$\cong\ker\Big(H^{4}_{|\widetilde{Z}_{0}|}(\widetilde{X}_{0};\mathds{Q}(2))\to H^{4}(\widetilde{X}_{0};\mathds{Q}(2))\Big)=\langle\{\widetilde{Z}_{0})\}\rangle.$
Also, it is evident that the unique generator $1_{t}\in\mathds{Q}(0)$ maps to $\{Z_{0}\}$, which in turn maps to $\{\widetilde{Z}_{0}\}$ under $\nu^{\ast}$. The subscript $t$ reflects the choice of coordinate made in assumption 2.2.
Now, invoking the functoriality of limit mixed Hodge structures once again, we know that the duality isomorphism $H^{3}(JC_{t}\setminus|W_{t}|;\mathds{Q}(2))^{\vee}\cong H^{3}(JC_{t},|W_{t}|;\mathds{Q}(2))$ carries over to the limit. Thus, we obtain the limit of the dual Abel-Jacobi image of $W_{t}$.
(4.7) $0\to\mathds{Q}(1)\to H^{3}_{\lim}(JC_{t},|W_{t}|;\mathds{Q}(2))\to H^{3}_{\lim}(JC_{t};\mathbb{Q}(2))\to 0.$
Reasoning as before for $\mathds{Q}(0)$, we obtain that the $\mathds{Q}(1)$ in (4.7) is generated by the dual of $\operatorname{cl}(W_{0})$ (or dual of $\operatorname{cl}(\widetilde{W}_{0})$ using $\nu$). Finally, combining equations (4.5) and (4.7), we obtain the limit MHS for the biextension variation $B_{Z_{t},W_{t}}=H^{3}(JC_{t}\smallsetminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$. Concretely, we get the limit version of diagram 1:
[…]
Recall the notation $\mathcal{V}_{t}=H^{3}(JC_{t}\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$. From the above diagram and the functoriality of mixed Hodge structures, it is easy to conclude that
$\operatorname{Gr}_{k}^{M}\mathcal{V}_{\lim}=\begin{cases}\mathds{Q}(0)\oplus H^{2}(J\widetilde{C}_{0};\mathds{Q}(1)),&\text{ if }k=0,\\ H^{3}(\widetilde{X}_{0};\mathds{Q}(2)),&\text{ if }k=-1,\\ \mathds{Q}(1)\oplus H^{2}(J\widetilde{C}_{0};\mathds{Q}(2)),&\text{ if }k=-2,\\ 0,&\text{ otherwise}.\end{cases}$
Let $\delta_{M}$ be the Deligne delta associated with this limit mixed Hodge structure. From the shape of the limit mixed Hodge structure, $\delta_{M}\colon\operatorname{Gr}^{M}_{0}\mathcal{V}_{\lim}\to\operatorname{Gr}^{M}_{-2}\mathcal{V}_{\lim}$ with respect to the grading $Y_{(F_{\infty},M)}$. Now, let $Y=Y(N,F_{\infty},M)$ be the unique grading of the original filtration $W$, as described in subsection 3.2. The component $\delta_{M,-2}$ modulo $\text{ad}(Y)$ is given by the component of $\delta_{M}$ that goes from $\mathds{R}(0)$ to $\mathds{R}(1)$ (cf. the discussion following Remark 3.8), while $\delta_{M,-1}=\delta^{A}_{M,-1}+\delta^{B}_{M,-1}$, where
$\delta^{A}_{M,-1}\colon\mathds{R}(0)\to H^{2}(J\widetilde{C}_{0};\mathds{R}(2)),$
$\delta^{B}_{M,-1}\colon H^{2}(J\widetilde{C}_{0};\mathds{R}(1))\to\mathds{R}(1).$
Finally, for analogous reasons, $\delta_{M,0}$ is given only by
$\delta_{M,0}\colon H^{2}(J\widetilde{C}_{0};\mathds{R}(1))\to H^{2}(J\widetilde{C}_{0};\mathds{R}(2)),$
which, by the choice of coordinate, is trivial.
By equation (3.2), $\delta_{M,-2}$ defines the limit height $H(N,F_{\infty},W)$. Next, let us recall that, by Remark 4.2, $\widetilde{Z}_{0},$ and $\widetilde{W}_{0}$, are rationally trivial in $Z^{2}(\widetilde{X}_{0})\otimes\mathds{Q}$. This implies that the Abel-Jacobi class $\operatorname{AJ}(\widetilde{Z}_{0})\in JH^{3}(\widetilde{X}_{0};\mathds{Q}(2))$ is the splitting extension class
(4.8) $0\to H^{3}(\widetilde{X}_{0},\mathds{Q}(2))\to H^{3}(\widetilde{X}_{0},\mathds{Q}(2))\oplus\mathds{Q}(0)\to\mathds{Q}(0)\to 0.$
The same holds for $\operatorname{AJ}(\widetilde{W}_{0})$ and its dual. Hence, diagram 1 simplifies, and the biextension MHS acquires the shape
$B_{\widetilde{Z}_{0},\widetilde{W}_{0}}=H^{3}(\widetilde{X}_{0}\smallsetminus|\widetilde{Z}_{0}|,|\widetilde{W}_{0}|;\mathds{Q}(2))\cong\mathds{Q}(0)\oplus H^{3}(\widetilde{X}_{0},\mathds{Q}(2))\oplus\mathds{Q}(1)$
as a $\mathds{Q}$-vector space, with only one non-trivial mixed Hodge component in $\operatorname{Ext}^{1}(\mathds{Q}(0),\mathds{Q}(1))\cong\mathds{C}/(2\pi i)\mathds{Q}$, and the corresponding height $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$.
By Proposition 4.1, the family of Ceresa cycles $Z_{t}$ extends to $Z_{0}$, which is homologous to zero in $H^{4}(X_{0};\mathds{Q}(2))$. Thus, the localization sequence together with Proposition 4.3 gives us the Abel Jacobi element $AJ(Z_{0})$ as the class of the extension
(4.9) $0\to H^{3}(X_{0};\mathds{Q}(2))\to H^{3}(X_{0}\smallsetminus|Z_{0}|;\mathds{Q}(2))\to\mathds{Q}(0)\to 0.$
Using the functoriality of the graded pieces with respect to morphisms of mixed Hodge structures, and (1.6), we get the graded pieces of $H^{3}(X_{0}\smallsetminus|Z_{0}|;\mathds{Q}(2))$:
$\operatorname{Gr}_{k}^{W}H^{3}(X_{0}\smallsetminus|Z_{0}|;\mathds{Q}(2))=\begin{cases}\mathds{Q}(0),&\text{ if }k=0,\\ H^{3}(\widetilde{X}_{0};\mathds{Q}(2)),&\text{ if }k=-1,\\ H^{2}(J\widetilde{C}_{0};\mathds{Q}(2)),&\text{ if }k=-2,\\ 0,&\text{ otherwise.}\end{cases}$
Similarly, $AJ(W_{0})$ is given by the class of the extension:
(4.10) $0\to H^{3}(X_{0};\mathds{Q}(2))\to H^{3}(X_{0}\smallsetminus|W_{0}|;\mathds{Q}(2))\to\mathds{Q}(0)\to 0.$
Using Verdier duality, we get (after twisting by $\mathds{Q}(1)$):
$0\to\mathds{Q}(1)\to H^{3}(X_{0},|W_{0}|;\mathds{Q}(2))\to H^{3}(X_{0};\mathds{Q}(2))\to 0.$
It follows that combining (4.9) with the dual of (4.10), we get the MHS $H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))$ and a diagram similar to diagram 1, with the mixed Hodge structure $H^{3}(X_{0};\mathds{Q}(2))$ on the top left corner. Hence, we get the graded pieces
$\operatorname{Gr}_{k}^{W}H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))=\begin{cases}\mathds{Q}(0),&\text{ if }k=0,\\ H^{3}(\widetilde{X}_{0};\mathds{Q}(2)),&\text{ if }k=-1,\\ \mathds{Q}(1)\oplus H^{2}(J\widetilde{C}_{0};\mathds{Q}(2)),&\text{ if }k=-2,\\ 0,&\text{ otherwise.}\end{cases}$
Using the relative version of Theorem 2.1, we have a morphism of mixed Hodge structures
$\iota:H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))\to H^{3}_{lim}(JC_{t}\smallsetminus|Z_{t}|,|W_{t}|;\mathds{Q}(2)).$
Indeed, the inclusion $H^{1}(C_{0};\mathds{Q})\hookrightarrow H^{1}_{\lim}(C_{t};\mathds{Q})$, and the fact that $H^{3}_{\lim}(JC_{t};\mathds{Q})=\wedge H^{1}_{\lim}(C_{t};\mathds{Q})$, shows that $H^{3}(X_{0};\mathds{Q}(2))\hookrightarrow H^{3}_{\lim}(JC_{t};\mathds{Q}(2))$. Further, using the five lemma, we obtain
$H^{3}(X_{0}\setminus|Z_{0}|;\mathds{Q}(2))\hookrightarrow H^{3}_{\lim}(JC_{t}\setminus|Z_{t}|;\mathds{Q}(2)),$
and using (relevant versions of) diagram 1, we conclude that $\iota$ is also injective. Let us denote $\mathcal{Q}=\operatorname{Im}(\iota)\subset\mathcal{V}_{\lim}$. We note that $\operatorname{Gr}_{k}\iota$ induces isomorphism for k=1,2k=-1,-2, while $\operatorname{Gr}_{0}\iota$ is injective. By (4.6), it is also evident that $\operatorname{Gr}_{0}\iota$ maps $\{Z_{0}\}$ to the canonical Betti generator $1_{t}\in\mathds{Q}(0)=\operatorname{Gr}^{M}_{0}\mathcal{V}_{\lim}$.
Now, let $\delta_{\mathcal{Q}}$ be the Deligne splitting associated with $\mathcal{Q}$ (which, using $\iota$, we identify with that of $H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))$). By Lemma 1.7, we have that
(4.11) $\delta_{M}|_{\mathcal{Q}}=\delta_{\mathcal{Q}}.$
Therefore, if we restrict to the component $\delta^{-1,-1}_{\mathcal{Q}}\colon\mathds{R}(0)\to\mathds{R}(1)$ of $\delta_{\mathcal{Q}}$ and to $\delta_{M,-2}$ of $\delta_{M}$, we read the condition (4.11), as $\delta^{-1,-1}_{\mathcal{Q}}=\delta_{M,-2}$. Hence, we can restate the limit height formula as
$\delta^{-1,-1}_{\mathcal{Q}}(1_{t})=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee}_{t},$
where, as before $1_{t}\in\mathds{Q}(0)$ and $1^{\vee}_{t}\in\mathds{Q}(1)$ are the canonical Betti generators. Now, define
$f\coloneqq\nu^{\ast}\circ\iota^{-1}\colon\mathcal{Q}\to B_{\widetilde{Z}_{0},\widetilde{W}_{0}},$
as a morphism of mixed Hodge structures. It is easy to see that $f(1_{t})=1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ and $f(1^{\vee}_{t})=1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$, where we recall that $1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by $\{\widetilde{Z}_{0}\}$, while $1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$ is given by the dual of $\{\widetilde{W}_{0}\}$. Since $f$ is a morphism of MHS, by Lemma 1.7 we have $f\circ\delta_{\mathcal{Q}}=\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Notice that since $B_{\widetilde{Z}_{0},\widetilde{W}_{0}}$ is a biextension MHS, we have $\delta_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}$. On the other hand, $\mathcal{Q}$ is not a biextension mixed Hodge structure. But from the shape of its graded components we can conclude once again, using Lemma 1.7, and strictness of $f$, that $f\circ\delta^{-1,-1}_{\mathcal{Q}}=\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}\circ f$. Now we can invoke the argument of Proposition 3.4 in this situation to conclude
$\delta^{-1,-1}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}(1_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}})=\frac{1}{2\pi}H(N,F_{\infty},W)1^{\vee}_{B_{\widetilde{Z}_{0},\widetilde{W}_{0}}}.$
As an immediate consequence, we get $H(N,F_{\infty},W)=\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$.
The case when $C_{0}=C_{1}\cup C_{2}$ is reducible follows the same set of ideas. In this scenario, we have $H^{1}_{\lim}(C_{t};\mathds{Q})\cong H^{1}(C_{1};\mathds{Q})\oplus H^{1}(C_{2};\mathds{Q})$, which is a pure Hodge structure and is automatically split. Hence, unlike the previous scenario, a choice of a coordinate cannot be made at this stage. In this case, the biextension variation $B_{Z_{t},W_{t}}=H^{3}(J(C_{t})\setminus|Z_{t}|,|W_{t}|;\mathds{Q}(2))$ extends to a biextension mixed Hodge structure $B_{Z_{0},W_{0}}=H^{3}(X_{0}\smallsetminus|Z_{0}|,|W_{0}|;\mathds{Q}(2))$, where $X_{0}=JC_{1}\times JC_{2}$, and $M=W$. Notice that what we really have is the following: For any coordinate $t\in\Delta$, the limit of the biextension variation $B_{Z_{t},W_{t}}$ is isomorphic to $B_{Z_{0},W_{0}}$. However, this isomorphism is not fixed and depends on $t$. In this case, one can choose a coordinate $t$ in such a way that (as before) $1_{t}\mapsto\{Z_{0}\}$. In any case, we have that $H(N,F,W)$ is given by the Deligne $\delta$ corresponding to $B_{Z_{0},W_{0}}$. Notice that, in this case, $Z_{0}$ and $W_{0}$ are homologous to zero, but may not be rationally equivalent to zero (Proposition 4.4). ∎
We end with a small remark.
Remark 4.8.
Even though the cycles $\widetilde{Z}_{0}$ and $\widetilde{W}_{0}$ are rationally equivalent to zero (after tensoring with $\mathds{Q}$), the height $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$ could be non-zero. In fact, we can get a formula for the height in terms of logarithms. Since $\widetilde{Z}_{0}$ is rationally equivalent to zero (up to torsion), we know that there exist finitely many divisors $Y_{i}$, and rational functions $f_{i}\in\mathds{C}(Y_{i})^{\ast}$, such that $\widetilde{Z}_{0}=\Sigma_{i}\operatorname{div}(f_{i})$. One can also arrange the $Y_{i}$-s to intersect $\widetilde{W}_{0}$ properly. Let $Y_{i}\cdot\widetilde{W}_{0}=\Sigma_{j}n_{ij}p_{ij}$, for each $Y_{i}$. Then, following Remark 3.6, one can show that $\operatorname{ht}(B_{\widetilde{Z}_{0},\widetilde{W}_{0}})$ is given by $\Sigma_{i,j}n_{ij}\log|f_{i}(p_{ij})|$.